OpenStudy (anonymous):

sales tax is 7% and the sales tax collected is $214.20. find the original price of the item being sold

OpenStudy (anonymous):

This question is really asking: "214.20 is 7% of what number ?" 0.07 x = 214.20 Can you solve that for x? That'll give you the answer.

OpenStudy (anonymous):

X * 0.93 = 214.20 X = 214.2 / 0.93 X = about 230.32

OpenStudy (anonymous):

yes bangokgarrett thank you!

OpenStudy (anonymous):

That's incorrect, @D35TR0Y3R .

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Ohhh, misread the question. Sorry

OpenStudy (anonymous):

Then X * 0.07 = 214.2 X = 214.2 / 0.07 X = $3060 I just put X on one side and then did simple arithmetic